Ncert Solutions Maths class 12th
Get insights from 2.5k questions on Ncert Solutions Maths class 12th, answered by students, alumni, and experts. You may also ask and answer any question you like about Ncert Solutions Maths class 12th
- 2.5k
Questions
- 0
Discussions
- 1
Active Users
- 65
Followers
Have a question about Ncert Solutions Maths class 12th? Ask our experts
Keep it short & simple. Type complete word. Avoid abusive language.
54. From a lot of 30 bulbs which include 6 defectives, a sample of 4 bulbs is drawn at random with replacement. Find the probability distribution of the number of defective bulbs.
54. From a lot of 30 bulbs which include 6 defectives, a sample of 4 bulbs is drawn at random with replacement. Find the probability distribution of the number of defective bulbs.
54. It is given that out of 30 bulbs, 6 are defective.
⇒ Number of non-defective bulbs = 30 − 6 = 24
4 bulbs are drawn from the lot with replacement.
Let X be the random variable that denotes the number of defective bulbs in the selected bulbs.
(4 non-defective and 0 defective)
(3 non-d
53. Find the probability distribution of the number of success in two tosses of a die where a success is defined as:
(i) Number greater than 4.
(ii) Six appears on at least one die.
53. Find the probability distribution of the number of success in two tosses of a die where a success is defined as:
(i) Number greater than 4.
(ii) Six appears on at least one die.
53. When a die is tossed two times, we obtain (6 × 6) = 36 number of observations.
Let X be the random variable, which represents the number of successes.
(i) Here, success refers to the number greater than 4.
P (X = 0) = P (number less than or equal to 4 on both the tosses) = 4/6 X 4/6 = 4/9
P (X = 1)
52. Find the probability distribution of:
(i) Number of heads in two tosses of a coin.
(ii) Number of tails in the simultaneous tosses of three coins.
(iii) Number of heads in four tosses of a coin.
52. Find the probability distribution of:
(i) Number of heads in two tosses of a coin.
(ii) Number of tails in the simultaneous tosses of three coins.
(iii) Number of heads in four tosses of a coin.
(i) When one coin is tossed twice, the sample space is
{HH, HT, TH, TT}
Let X represent the number of heads.
∴ X (HH) = 2, X (HT) = 1, X (TH) = 1, X (TT) = 0
Therefore, X can take the value of 0, 1, or 2.
It is known that,
P (HH) = P (HT) = P (TH) = P (TT) = 1/4
P (X = 0) = P (TT) = 1/4
P (X = 1) =
51. Let X represents the difference between the number of heads and the number of tails obtained when a coin is tossed 6 times. What are possible values of X?
51. Let X represents the difference between the number of heads and the number of tails obtained when a coin is tossed 6 times. What are possible values of X?
51. A coin is tossed six times and X represents the difference between the number of heads and the number of tails.
∴ X (6 H, 0T) = |6 - 0| = 6
X (5 H, 1 T) = |5 - 1| = 4
X (4 H, 2 T) = |4 - 2| = 2
X (3 H, 3 T) = 3 - 3| = 0
X (2 H, 4 T) = |2 - 4| = 2
X (1 H, 5 T) =
50. An urn contains 5 red and 2 black balls. Two balls are randomly selected. Let X represents the number of black balls. What are the possible values of X? Is X a random variable?
50. An urn contains 5 red and 2 black balls. Two balls are randomly selected. Let X represents the number of black balls. What are the possible values of X? Is X a random variable?
50. There two balls may be selected as BR, RB, BR, BB, where R represents red ball and B represents black ball.
Variable X has the value 0, 1, 2, i.e., there may be no black ball, may be one black ball or both the balls are black.
49. State which of the following are not the probability distributions of a random variable. Give reasons for your answer.
(i)
X
0
1
2
P (X)
0.4
0.4
0.2
(ii)
X
0
1
2
3
4
P (X)
0.1
0.5
0.2
– 0.1
0.3
(iii)
Y
– 1
0
1
P (Y)
0.6
0.1
0.2
(iv)
Y
– 1
0
1
P (Y)
0.6
0.1
0.2
49. State which of the following are not the probability distributions of a random variable. Give reasons for your answer.
(i)
|
X |
0 |
1 |
2 |
|
P (X) |
0.4 |
0.4 |
0.2 |
(ii)
|
X |
0 |
1 |
2 |
3 |
4 |
|
P (X) |
0.1 |
0.5 |
0.2 |
– 0.1 |
0.3 |
(iii)
|
Y |
– 1 |
0 |
1 |
|
P (Y) |
0.6 |
0.1 |
0.2 |
(iv)
|
Y |
– 1 |
0 |
1 |
|
P (Y) |
0.6 |
0.1 |
0.2 |
49. It is known that the sum of all the probabilities in a probability distribution is one.
(i) Sum of the probabilities = 0.4 + 0.4 + 0.2 = 1
Therefore, the given table is a probability distribution of random variables.
(ii) It can be seen that for X = 3, P (X) = −0.1
It is known that proba
48. If A and B are two events such that A ⊂ B and P (B) ≠ 0, then which of the following is correct:
48. If A and B are two events such that A ⊂ B and P (B) ≠ 0, then which of the following is correct:
48. If then
also
Consider,
We know,
From eq. , we have
is not less than
Hence, from eq. it can be concluded that the relation given in alternative is correct.
47. Probability that A speaks truth is 4/5 A coin is tossed. A report that a head appears. The probability that actually there was head is:
(A) 4/5
(B) 1/2
(C) 1/5
(D) 2/5
47. Probability that A speaks truth is 4/5 A coin is tossed. A report that a head appears. The probability that actually there was head is:
(A) 4/5
(B) 1/2
(C) 1/5
(D) 2/5
47. Let, and be the event such that
speak truth
speak false
that head appears
and =
If a coin tossed, then it may result either head or tail . The probability of getting a head is whether &nbs
46. A card from a pack of 52 cards is lost. From the remaining cards of the pack, two cards are drawn and are found to be both diamonds. Find the probability of the lost card being a diamond.
46. A card from a pack of 52 cards is lost. From the remaining cards of the pack, two cards are drawn and are found to be both diamonds. Find the probability of the lost card being a diamond.
46. Let, event of choosing a diamond card
event of choosing a card which is not diamond
denote the lost card, out of cards
We know,
cards are diamond
cards are not diamond
When one diamond card is lost, there are diamond cards out of card
45. A manufacturer has three machine operators A, B and C. The first operator A produces 1% defective items, whereas the other two operators B and C produce 5% and 7% defective items respectively. A is on the job for 50% of the time, B on the job for 30% of the time and C on the job for 20% of the time. A defective item is produced, what is the probability that it was produced by A?
45. A manufacturer has three machine operators A, B and C. The first operator A produces 1% defective items, whereas the other two operators B and C produce 5% and 7% defective items respectively. A is on the job for 50% of the time, B on the job for 30% of the time and C on the job for 20% of the time. A defective item is produced, what is the probability that it was produced by A?
45. Let, event of time consume by
event of time consume by
event of time consume by
Let, event of producing the defective item. Therefore,
Therefore, by Baye’s theorem,
probability that the defective item was produced by ,